HIGH RESOLUTION ELECTRONIC SPECTROSCOPY OF METHYL ANISOLES.^{a}

\underline{LEONARDO ALVAREZ}, JOHN T. YI and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

Rotationally resolved S_{1} ← S_{0} fluorescence excitation spectra of 2-methyl anisole (2MA) and 3-methyl anisole (3MA) have been obtained in the environment of a supersonic jet environment. The derived values of the rotational constants show that 2MA itself has a planar heavy-atom structure with trans-disposed methoxy groups in both electronic states. No evidence for any other conformational isomers was found. Also, 3MA itself has a planar heavy-atom structure but with trans- and cis-disposed methoxy groups in both electronic states. Even though both molecules exhibit hindered internal methyl rotation, the spectrum of 2MA shows only a single band, while 3MA shows multiple sub-bands within each band. These spectral features will be discussed further in detail.

aWork supported by NSF.